Sol Summary – January 20th

MDRS Sol 5 Summary
Summary Title: Hot Snow and Cool Cats
By GreenHab Officer Richard Blake
Mission Status:
Captain‘s log:: Stardate 002017J1 – We have … encountered some kind of … frozen liquid. It coats everything … crunches beneath your boots. It may … hamper the mission, but we will … push on ::end log.
Sol Activity Summary:
Dawn’s rays illuminated a pristine white landscape, draped in a thin blanket of snow. It was the first time we have seen a snow covered Martian landscape and gave us the perfect opportunity to break out the cameras and attempt to capture the beauty of our natural setting.
Michaela and Idriss went on an EVA (described below) to attempt to collect some extremophile samples, despite the snowy and muddy conditions. They were successful in their mission and they also succeeded in driver the rover in these difficult conditions, which was something the crew has attempted for the first time.
When Michaela and Idriss got back to the Hab from their EVA, the whole crew donned their exo suits and exited the airlock to join them. The aim of this mass EVA was to take photgraphic advantage of the snowy conditions. Niamh (aka Midas) worked her magic and captured a number of stunning shots of both the team and the landscape.
Around midday a pair of bipedal aliens parked their speeder a short distance from the Hab. They quickly fortified a position inside our GreenHab, spending a good few hours studying our technology. When they retreated, we found to our delight that they had somehow improved our life support systems, leaving the GreenHab capable of supporting our plants overnight!
As the sun shone down around us in the afternoon, it turned the beautiful white snow to a muddy slush, leaving the ground as muddy as it was when we first arrived. This will slow any EVA progress in the near future, but hopefully a few sunny days will dry it out again.
Idriss has been hard at work trying to get the 3D printer up and running again. Hopefully his efforts will be rewarded and his goals of printing bricks achieved in the coming days (if not hours!).
Roy has also taken over the Science Dome, using the facilities to begin studying the many rock and soil samples he has collected over the past week. A number of interesting features were found, including a fascinating iron nodule in one sample.
Michaela is continuing to attend to her students‘ project and her projects in collaboration with the Czech Masaryk and Mendel Universities. I will be helping her out on these, as they involve a lot of GreenHab related activities.
Tonight will conclude with Irish culture night, where we are sure to eat many a potato, joke about leprechauns and cry that we can’t sate ourselves with a nice Guinness.
Look Ahead Plan:
With the GreenHab heater fully functional now, more seedlings can be grown, transplanted and left in there permentantly.
The geology lab-work can also begin in earnest with Roy and myself logging and recording data on the samples collected. I will also continue helping Michaela with her numerous projects.
Depending on weather and ground conditions, an EVA will be conducted by Roy to collect more samples relevant for his research and Michaela will try to capture a few videos/photos for outreach purposes.
Idriss will be continuing his 3D printing work, and either still be troubleshooting or printing bricks like Pokemon Go prints money.
Anomalies in work:
Aliens took over the GreenHab for a few hours, however they seem only to have upgraded our life support systems in there. No other anomalies to report.
Landscape was snow covered in the morning, and overcast. It warmed up a little in the afternoon, and the sun even poked his head out for an hour or so. Snow has mostly melted now.
Crew Physical Status:
Tip top – though I suspect the isolation is getting to a certain crew-mate. We may not be safe for much longer… 😉
An EVA was conducted in the morning, when the ground was still not too muddy. The purpose of this was to collect some hypolith samples for Michaela. The snow had melted enough to allow for a few promising samples to be collected.
Reports to be filed:
Sol 5 EVA report
Sol 6 EVA request
Journalist’s report
Operations report
Support Requested:
None today